# Efficient and Responsible Adaptation of Large Language Models for Robust Top-k Recommendations

### v2

@@ 5. Discussion
This paper presents a foundational framework for the responsible adaptation of Large Language Models (LLMs) in ranking tasks, with particular attention to the under-explored challenge of handling weak users. Our findings reveal that traditional recommendation systems (RS) exhibit significant weaknesses in serving these marginalized users. The empirical evidence (see Fig. 7) shows that many weak users receive poor recommendations, often because the systems struggle to capture their latent preferences due to limited data. However, some weak users receive comparable recommendations to more active users, which suggests that traditional RS can still identify certain common patterns based on item similarity. We also leveraging in-context learning to contextualize the sparse rating histories of these weak users, adapting the model to recognize nuanced preferences that traditional RSs miss out. This shift from group-based evaluation metrics to an instance-by-instance approach directly aligns with recent calls in the literature, particularly from Burnell et al. (2023), emphasizing the need for tailored, high-stakes decision-making systems.

### v3

@@ 5. Discussion
This paper presents a foundational framework for the responsible adaptation of Large Language Models (LLMs) in ranking tasks, with particular attention to the under-explored challenge of handling weak users. Our findings reveal that traditional recommendation systems (RS) exhibit significant weaknesses in serving these high-risk users. The empirical evidence (see Fig. 7) shows that many weak users receive poor recommendations, often because the systems struggle to capture their latent preferences due to limited data. However, some weak users receive comparable recommendations to more active users, which suggests that traditional RS can still identify certain common patterns based on similarity in rating patterns. We also leverage in-context learning to contextualize the sparse rating histories of these weak users, adapting the model to recognize nuanced preferences that traditional RSs miss out. This shift from group-based evaluation metrics to an instance-by-instance approach directly aligns with recent calls in the literature, particularly from Burnell et al. (2023), emphasizing the need for tailored, high-stakes decision-making systems.
While Table 3 demonstrates that multiple LLM-augmented models improve performance, GPT-4 consistently outperforms other LLMs across weak and non-weak user groups. We attribute this to GPT-4’s superior in-context learning capabilities, enabling it to generalize effectively from sparse and noisy user histories. Its larger parameter count and deeper pretraining allow it to interpret vague or limited preferences better than smaller LLMs like LLaMA or Claude. Additionally, GPT-4 appears more robust in recognizing item semantics and aligning them with implicit user intent, particularly in cold-start scenarios. This observation supports our motivation for selective LLM invocation: while GPT-4 offers strong gains, especially for weak users, its computational cost justifies using it only when traditional RS models underperform. These findings reinforce our framework’s core hypothesis—that task-aware allocation of LLM resources can deliver both performance improvements and efficiency.